find the length of a rectangular lot with the perimeter of 132 meters if the length is 8 meters more than the width
Answer by mananth(16949)

let width be x
length =x+8
..
Perimeter = 2*L+2*W
2x+2(x+8)=132
2x+2x+16=132
4x+16=132
4x=116
x=29 m which is the width
the length = 29+8=37 m
..